7  Kim's Fur Shoppe has 11,200 shares of common stock outstanding at a price of $46 per share. It also has 265 shares of preferred stock outstanding at a price of $88 per share. There are 620 bonds outstanding that have a coupon rate of 6.5 percent paid semiannually. The bonds mature in 27 years, have a face value of $1,000, and sell at 107 percent of par. What is the capital structure weight (WE) of the common stock?

A               0.5018                    B               0.447                       C               0.4286                    D               0.3256
Answer #1
MV of equity=Price of equity*number of shares outstanding
MV of equity=46*11200
=515200
MV of Bond=Par value*bonds outstanding*%age of par
MV of Bond=1000*620*1.07
=663400
MV of Preferred equity=Price*number of shares outstanding
MV of Preferred equity=88*265
=23320
MV of firm = MV of Equity + MV of Bond+ MV of Preferred equity
=515200+663400+23320
=1201920
Weight of equity = MV of Equity/MV of firm
Weight of equity = 515200/1201920
W(E)=0.4286